OpenStudy (anonymous):

A spring stretches 0.05 m when a 0.3 kg mass is suspended from it. What is the characteristic frequency of this mass on the spring?

OpenStudy (anonymous):

i get : \[1/2\pi \sqrt{k/m} = 2.228 Hz\] does anyone know if its right?